a conducting cylinder of radius 3.25 cm and length 5.6 cm has a total charge of 4.5 x 10 c distributed uniformly on its surface area. find the potential at a) its surface, and b) 2.5 cm and c) 5.0 cm from the center of the cylinder.

1
Expert's answer
2022-02-21T12:05:45-0500

The linear density of charge

"\lambda=\frac{Q}{l}=\frac{45\:\rm C}{0.0325\:\rm m}=45\:\rm C/m"

The potential of electric field from charged cylinder

"V=2k\lambda\ln\frac{r}{r_0}"

(a) "r=r_0"

"V=0"

(b) "r=2.5\:\rm cm"

Inside a conducting cylinder the field is constant and equal to the it's surface value. Hence

"V=0"

(c) "r=5.0\:\rm cm"

"V=2*9*10^9*45*\ln\frac{5.0}{3.25}=152*10^9\;\rm V"